Return question 9 record declared a $1.50 per bare cash dividend payable on February 28 to the February 9 stockholders of AleReturn to question 9 1 Jan 02 125,000 Treasury stock, common Cash 125,000 2 Jan 07 67,500 1.12 points Retained eamings Common50,000 9 Treasury stock, common Paid-in capital, treasury stock 10,000 5 Aug 27 Cash 1.12 points Paid-in capital, treasury st9 Complete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 1.12 points Prepare a sComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 1.12 points Prepare the sI tried every number, and could not find the correct answer for Sep. 9!

Please help!!
Answer #1
1
Date General Journal Debit Credit
2-Jan Treasury stock, common 125,000 =5000*25
Cash 125,000
7-Jan Retained earnings 67,500 =(50000-5000)*1.5
Common dividend payable 67,500
28-Feb Common dividend payable 67,500
Cash 67,500
9-Jul Cash 60,000 =2000*30
Treasury stock, common 50,000 =2000*25
Paid-in capital, treasury stock 10,000
27-Aug Cash 50,000 =2500*20
Paid-in capital, treasury stock 10,000
Retained earnings 2,500
Treasury stock, common 62,500 =2500*25
9-Sep Retained earnings 99,000 =(50000-5000+2000+2500)*2
Common dividend payable 99,000
22-Oct Common dividend payable 99,000
Cash 99,000
31-Dec Income summary 72,000
Retained earnings 72,000
2
ALEXANDER CORPORATION
Statement of Retained Earnings
For Year Ended December 31, 2017
Retained earnings, December 31, 2016 410,000
Add: Net income 72,000
482,000
Less: Cash dividends declared -166,500
Less: Treasury stock reissuances -2,500
Retained earnings, December 31, 2017 313000
3
ALEXANDER CORPORATION
Stockholders’ Equity Section of the Balance Sheet
December 31, 2017
Common stock - $25 par value 1,250,000
Paid-in capital in excess of par value, common stock 100,000
Retained earnings 313,000
Less cost of treasury stock -12,500
Total stockholders’ equity 1,650,500